what are the frequency and wavelength of a photon emitted during a transition from n=5 state to then n=2 state in the hydrogen atom

Grade:11

1 Answers

Vikas TU
7 years ago
From wave number formulae:
1/k = Rz^2*(1/2^2 – 1/5^2)
1/k = R*1^2*(1/4 – 1/25)
1/k = 21R/100
k = 100/21R
f = c/k put he k value.